# Envelopes in the class of Banach algebras of polynomial growth and $C^{\infty}$-functions of a finite number of free variables

### 11 Jan 2024 v1

@@ 5. Envelopes of free algebras
When $\mathsf{C}=\mathsf{PG}$, tensor $\mathsf{CL}$ algebras can be naturally called tensor $\mathsf{PGL}$ algebras and, particularly, free $\mathsf{CL}$ algebras can be called free $\mathsf{PGL}$ algebras. We also name them $C^{\infty}$-tensor algebras and $C^{\infty}$-free algebras. In this section we describe an explicit form of the free $\mathsf{PGL}$ algebra $\mathscr{F}^{\,\mathsf{PG}}\{X\}$ in the case when $X$ is a finite set of cardinality $k$. For brevity, we denote $\mathscr{F}^{\,\mathsf{PG}}\{X\}$ by $\mathscr{F}_{k}^{\,\mathsf{PG}}$. The proof is given only for $k\leqslant 2$; the general case is postponed to [8].
@@
For a proof (in the case of a free Lie algebra over an arbitrary commutative ring) see [9, p. 64, § 2.4.2, Corollary 2.16(ii)] or [13, p. 72, Lemma 2.11.16].
@@
where $a$ and $b$ are elements of a Banach algebra and $s$ is a scalar; see, e.g., [11, Chapter II, § 15, p. 83, Remark 1].
@@
is absolutely integrable. Thus, the integral exists not only in the weak sense (of Gelfand and Pettis), but also in the strong sense (of Bochner) [22, p. 80, Theorem 3.7.4]. Applying Theorem 4.5 again, now to the tuple $(\mathop{\mathrm{ad}}\nolimits b_{1},\ldots,\mathop{\mathrm{ad}}\nolimits b_{k})$, we see that $\Psi_{\delta}$ is indeed an extension of the ordered calculus in (5.4). From the definition of $C^{\infty}(X)$ (see (5.5)), we have an extension of the map $V\to B$ to $C^{\infty}(X)\to B$.
@@
These results show that $C^{\infty}_{\mathfrak{f}_{k}}$ are deserved to be named algebras of ‘free $C^{\infty}$- functions’. The proof of Theorem 5.6 is quite technical. We give a proof in the case when $k\leqslant 2$, which is a little simpler. The case of general $k$ will be published in a subsequent paper [8].
@@
When $k=1$, the set $X$ is empty and then $C^{\infty}_{\mathfrak{f}_{1}}=C^{\infty}(\mathbb{R})$. So the result is implied by the corresponding assertion in the commutative case, Proposition 3.3 with $k=1$.
@@
Let $m\in\mathbb{N}$ and $w=(\mathbf{x}_{1},\ldots,\mathbf{x}_{m})\in X^{m}$, where $x_{p}=(s_{p},t_{p})$. Consider the diagonal the following matrices:
@@
where $t_{p}^{\prime}=t_{p}+\cdots+t_{m}$, $s_{p}^{\prime}=s_{1}+\cdots+s_{p}$. Take the Jordan block $J$ of order $m+1$ with zero eigenvalue, i.e.,
@@
where $\mathfrak{c}_{2}$ denotes the set of matrices with zeros on two diagonals.
@@
It follows from (5.9), (5.10) and (5.14) that
$\theta_{w}(g_{\emptyset})=[F_{w}-J,E_{w}]=\begin{pmatrix}0&s_{1}&0&&\\ &0&s_{2}&\ddots&\\ &&\ddots&\ddots&0\\ &&&0&s_{m}\\ &&&&0\end{pmatrix}.$
@@
$Q_{m}\!:=\begin{pmatrix}0&0&\cdots&&1\\ &0&0&&\\ &&\ddots&\ddots&\vdots\\ &&&0&0\\ &&&&0\end{pmatrix}.$ (5.16)
@@
Note that the product of $m$ strictly upper-diagonal matrices may have non-zero entries everywhere except possibly the upper-right corner. This entry depends only on the entries on the second diagonals and equals the product of the $p$th entries in the $p$th matrix. Then it follows from Lemma 5.8 that (5.17) holds. ∎
@@
$\widetilde{\theta}_{\boldsymbol{\lambda},w}(\Phi_{m}(h))=s_{1}\cdots s_{m}\,h(\lambda_{1},\lambda_{2}+t_{1}+\cdots+t_{m},s_{1},t_{1},\ldots,s_{m},t_{m})Q_{m}.$ (5.19)
When $m=0$, the equality has the form $\widetilde{\theta}_{\boldsymbol{\lambda},\emptyset}(\Phi_{0}(h))=h(\lambda_{1},\lambda_{2})$.
@@
It is easy to see from the definitions that
@@
It follows from (5.4) and (5.18) that $\Phi_{m^{\prime}}(h)$ is a sum of products $m^{\prime}+1$ factors, where all the factors except first is of the form $g_{\beta}$, $\beta\in\mathbb{Z}_{+}^{2}$. Lemma 5.8 implies that $\theta_{w}(g_{\beta^{1}})\cdots\theta_{w}(g_{\beta^{m^{\prime}}})=0$ when $m<m^{\prime}$ and this completes the proof. ∎
In the proof of Theorem 5.6 we also need another lemma. First note that a family $(\psi_{ij}\!:X_{i}\to Y_{j};\,i\in I,\,j\in J)$ of continuous linear maps of complete locally convex spaces induces a continuous linear map $\prod_{i\in I}X_{i}\to\prod_{j\in J}Y_{j}$ under the assumption that for every $j$ the subset of $i$ such that $\psi_{ij}\neq 0$ is finite. Indeed, the finiteness condition implies that there is a morphism from $\prod_{i\in I}X_{i}$ to $Y_{j}$ for every $j$ and hence from $\prod_{i\in I}X_{i}$ to $\prod_{j\in J}Y_{j}$.
@@
It follows from a theorem of Dugundji [18, Theorem 4.1] that a continuous linear map between Fréchet spaces is topologically injective if and only if it has a continuous left inverse map (possibly, non-linear).
@@
Since $\prod B_{m}$ is in $\mathsf{PGL}$ by Corollary 2.8, it follows from Theorem 5.4 that $\rho$ extends to a continuous linear map from $C^{\infty}_{\mathfrak{f}_{2}}\to\prod B_{m}$. To complete the proof it suffices to show that $\rho$ is topologically injective.
Consider $B_{m}$ as a free $C^{\infty}(\mathbb{R}^{2}\times X^{m})$-module and denote by $Y_{m}$ the direct summand that corresponds to the upper-right corner (see the definition of the matrix $Q_{m}$ in (5.16)). It is easy to see that $\prod Y_{m}$ is a direct factor in $\prod B_{m}$. So, by Part (B) of Lemma 4.4, it suffices to show that the composition $\rho^{\prime}$ of $\rho$ and the projection on $\prod Y_{m}$ is topologically injective.
@@
The existence of an extension to an ordered $C^{\infty}$-calculus given by Theorem 5.4 implies that (5.19) in Lemma 5.10 holds for every $h\in C_{m}$. We claim that the endomorphism of $C^{\infty}(\mathbb{R}^{2}\times X^{m})$ given by

### 11 Mar 2024 v2

@@ 5. Envelopes of free algebras
When $\mathsf{C}=\mathsf{PG}$, tensor $\mathsf{CL}$ algebras can be naturally called tensor $\mathsf{PGL}$ algebras and, particularly, free $\mathsf{CL}$ algebras can be called free $\mathsf{PGL}$ algebras. We also name them $C^{\infty}$-tensor algebras and $C^{\infty}$-free algebras. In this section we describe an explicit form of the free $\mathsf{PGL}$ algebra $\mathscr{F}^{\,\mathsf{PG}}\{X\}$ in the case when $X$ is a finite set of cardinality $k$. For brevity, we denote $\mathscr{F}^{\,\mathsf{PG}}\{X\}$ by $\mathscr{F}_{k}^{\,\mathsf{PG}}$. The proof is given only for $k\leqslant 2$; the general case is postponed to a forthcoming paper.
@@
For a proof (in the case of a free Lie algebra over an arbitrary commutative ring) see [8, p. 64, § 2.4.2, Corollary 2.16(ii)] or [12, p. 72, Lemma 2.11.16].
@@
where $a$ and $b$ are elements of a Banach algebra and $s$ is a scalar; see, e.g., [10, Chapter II, § 15, p. 83, Remark 1].
@@
is absolutely integrable. Thus, the integral exists not only in the weak sense (of Gelfand and Pettis), but also in the strong sense (of Bochner) [21, p. 80, Theorem 3.7.4]. Applying Theorem 4.5 again, now to the tuple $(\mathop{\mathrm{ad}}\nolimits b_{1},\ldots,\mathop{\mathrm{ad}}\nolimits b_{k})$, we see that $\Psi_{\delta}$ is indeed an extension of the ordered calculus in (5.4). From the definition of $C^{\infty}(X)$ (see (5.5)), we have an extension of the map $V\to B$ to $C^{\infty}(X)\to B$.
@@
These results show that $C^{\infty}_{\mathfrak{f}_{k}}$ are deserved to be named algebras of ‘free $C^{\infty}$- functions’. The proof of Theorem 5.6 is quite technical. We give a proof in the case when $k\leqslant 2$, which is a little simpler. The case of general $k$ will be published in a subsequent paper.
@@
When $k=1$, the set $X$ is empty and then $C^{\infty}_{\mathfrak{f}_{1}}=C^{\infty}(\mathbb{R})$. So the result is implied by the corresponding assertion in the commutative case, Proposition 3.3 with $n=1$.
@@
Let $m\in\mathbb{N}$ and $w=(\mathbf{x}_{1},\ldots,\mathbf{x}_{m})\in X^{m}$, where $x_{p}=(s_{p},t_{p})$. For $m\geqslant 1$ consider the following diagonal matrices:
@@
where $t_{p}^{\prime}=t_{p}+\cdots+t_{m}$ and $s_{p}^{\prime}=s_{1}+\cdots+s_{p}$. For $m=0$ put $E_{\emptyset}=F_{\emptyset}=0$. Take the Jordan block $J$ of order $m+1$ with zero eigenvalue, i.e.,
@@
where $\mathfrak{c}_{2}$ denotes the set of triangular matrices with zeros on the first and second diagonals.
@@
It follows from (5.11) and (5.14) that
$\theta_{w}(g_{0})=[F_{w}-J,E_{w}]=\begin{pmatrix}0&s_{1}&0&&\\ &0&s_{2}&\ddots&\\ &&\ddots&\ddots&0\\ &&&0&s_{m}\\ &&&&0\end{pmatrix}.$
@@
$Q_{m}\!:=\begin{pmatrix}0&0&\cdots&&1\\ &0&0&&\\ &&\ddots&\ddots&\vdots\\ &&&0&0\\ &&&&0\end{pmatrix}$ (5.16)
when $m\geqslant 0$ and $Q_{0}=0$ (a scalar).
@@
Note that the product of $m$ strictly upper-diagonal matrices have zero entries everywhere except possibly the upper-right corner. This entry depends only on the entries on the second diagonals and equals the product of the $p$th entries in the $p$th matrix, where $p$ runs $\{1,\ldots,m\}$. Then it follows from Lemma 5.8 that (5.17) holds. ∎
@@
$\widetilde{\theta}_{\boldsymbol{\lambda},w}(\Phi_{m}(h))=s_{1}\cdots s_{m}\,h(\lambda_{1},\lambda_{2}+t_{1}+\cdots+t_{m},s_{1},t_{1},\ldots,s_{m},t_{m})Q_{m}$ (5.19)
when $m\geqslant 1$ and $\widetilde{\theta}_{\boldsymbol{\lambda},\emptyset}(\Phi_{0}(h))=h(\lambda_{1},\lambda_{2})$ when $m=0$.
@@
It is easy to see that
@@
It follows from (5.4) and (5.18) that $\Phi_{m^{\prime}}(h)$ is a sum of products of $m^{\prime}+1$ factors, where all the factors except first is of the form $g_{\beta}$, $\beta\in\mathbb{Z}_{+}^{2}$. Lemma 5.8 implies that $\theta_{w}(g_{\beta^{1}})\cdots\theta_{w}(g_{\beta^{m^{\prime}}})=0$ when $m<m^{\prime}$ and this completes the proof. ∎
In the proof of Theorem 5.6 we also need another lemma on topologically injective maps. First note that a family $(\psi_{ij}\!:X_{i}\to Y_{j};\,i\in I,\,j\in J)$ of continuous linear maps of complete locally convex spaces induces a continuous linear map $\prod_{i\in I}X_{i}\to\prod_{j\in J}Y_{j}$ under the assumption that for every $j$ the subset of $i$ such that $\psi_{ij}\neq 0$ is finite. Indeed, the finiteness condition implies that there is a morphism from $\prod_{i\in I}X_{i}$ to $Y_{j}$ for every $j$ and hence from $\prod_{i\in I}X_{i}$ to $\prod_{j\in J}Y_{j}$.
@@
It follows from a theorem of Dugundji [17, Theorem 4.1] that a continuous linear map between Fréchet spaces is topologically injective if and only if it has a continuous left inverse map (possibly, non-linear).
@@
Since $\prod B_{m}$ is in $\mathsf{PGL}$ by Corollary 2.8, it follows from Theorem 5.4 that $\rho$ extends to a continuous linear map from $C^{\infty}_{\mathfrak{f}_{2}}$ to $\prod B_{m}$. To complete the proof it suffices to show that $\rho$ is topologically injective.
Consider $B_{m}$ as a free $C^{\infty}(\mathbb{R}^{2}\times X^{m})$-module and denote by $Y_{m}$ the direct summand that corresponds to the upper-right corner (see the definition of the matrix $Q_{m}$ in (5.16)). It is easy to see that $\prod Y_{m}$ is a direct factor in $\prod B_{m}$. So, by Part (B) of Lemma 4.4, it suffices to show that the composition of $\rho$ and the projection on $\prod Y_{m}$ is topologically injective. Denote this composition by $\rho^{\prime}$.
@@
The existence of an extension to an ordered $C^{\infty}$-calculus given by Theorem 5.4 implies that (5.19) in Lemma 5.10 holds for every $h\in C_{m}$. We claim that the endomorphism of $C^{\infty}(\mathbb{R}^{2}\times X^{m})$ defined by